# A systematic review of challenges and proposed solutions in modeling multimodal data

### v2

@@ 1 Introduction
In healthcare, data comes in various modalities like electronic health records (EHR), imaging, genomics, and wearable devices. Integrating more than one of these modalities in a multimodal data analysis holds significant promise for improving our understanding and management of complex health conditions 1. One notable domain where multimodal data analysis can make a transformative impact is the management of chronic diseases like asthma and chronic obstructive pulmonary disease (COPD) patients 2. Here, medical imaging data, such as computerized tomography, provides insights into lung morphology and pathological changes. Wearable devices monitor physiological parameters and activity levels, offering real-time data on the patient’s disease trajectory. Environmental data, encompassing air quality and pollutants, helps identify external factors influencing respiratory conditions. Clinical records and genomic information contribute to personalized treatment strategies, while patient-reported outcomes capture subjective experiences, shaping a patient-centric approach. Combining several or even all of these modalities allows us to understand the disease and reveal interactions between different aspects.
@@
Modeling multimodal data is a challenging task that extends well beyond the analysis of individual modalities. Specifically, each modality may have different data types, formats, scales, levels of noise, and levels of interpretability. Moreover, modalities may exhibit varying degrees of completeness, sparsity, and dimensionality, making it challenging to integrate and analyze them together. In context of modeling multimodal data, fusion techniques are of great importance. A fusion strategy involves combining and integrating the different modalities to create a unified representation that captures the complementary information from each modality. These techniques can be divided into four main categories: (i) early fusion, (ii) intermediate fusion, (iii) late fusion, and (iv) hybrid fusion. (i) Early fusion involves integrating information from different sources at the data level without learning any marginal representations or feature extraction, treating them as unimodal input. Although early fusion offers simplicity and the ability to learn cross-modal relationships, it may pose challenges when dealing with different dimensionalities and sampling rates across modalities. Moreover, it overlooks the relationships between different modalities and may not be suitable for heterogeneous modalities, such as combining image data with omics data. (ii) Intermediate fusion, in contrast, involves learning modality-specific representations to identify correlations within each modality before generating joint representations or making predictions. This approach offers flexibility in selecting different architectures for feature extraction from various modalities 3. By employing distinct architectures and layers for each modality, it can effectively handle heterogeneous data. Additionally, intermediate fusion is resilient to dimensionality imbalances 3 and can accommodate missing modalities without significantly affecting performance 3, 4. (iii) Late fusion, on the other hand, trains separate models for each modality and integrates their individual predictions at a later stage. This method is particularly well-suited for managing heterogeneous modalities and addressing dimensionality imbalances and missing data. However, it may be less effective for modalities with strong correlations, as it does not capture interactions between different data sources. In addition, in scenarios where specific information holds greater significance, using late fusion offers notable benefits. The model prioritizes the most important information it has, then adds more details from other sources to improve its understanding. This prioritization allows for a more nuanced integration of information, enhancing the model’s ability to make informed decisions and predictions. Additionally, in scenarios where certain information is more critical, late fusion provides significant advantages. The model first prioritizes the most relevant information, then incorporates additional details from other sources to refine its understanding. This approach enables a more nuanced integration of data, improving the model’s ability to make informed decisions and predictions. (iv) Hybrid fusion combines elements from the aforementioned methods, allowing for flexible processing across multiple stages. This approach is especially effective in scenarios with diverse data modalities, where early fusion can be used for closely related data types, and later fusion stages facilitate the integration of more distinct modalities. While less frequently applied than the three primary fusion categories, hybrid fusion provides the adaptability needed to meet the specific requirements of multimodal models. However, selecting the optimal fusion technique represents only one of the many challenges faced in multimodal modeling 5. The integration of diverse data modalities into a multimodal model is visually represented in Figure 1. The figure visually represents the three primary fusion strategies (early, intermediate, and late), for simplicity, it excludes hybrid and other less frequently used strategies.
@@
The objective of this systematic review is to identify the challenges encountered in modeling multimodal data in medical domains. We aim to provide insight into the key challenges and their underlying causes, as well as potential solutions to address these challenges.
@@
Several review articles have discussed the application of deep learning to analyze multimodal data. For instance, Mohsen et al. 6 explored integrating multiple data modalities with artificial intelligence models for a variety of applications, including combining medical imaging with data from electronic health records (EHRs). Similarly, Azam et al. 7 and Zhang et al. 8 solely focused on fusing different imaging modalities. Behrad and Abadeh 9 and Stahlschmidt et al. 3 primarily concentrated on integrating omics data with other modalities. Huang et al. 10 address the fusion of multimodal EHR data using traditional machine learning and deep learning methods. Pawłowski et al. 11 and Shaik et al. 12 focus on fusion techniques and feature extraction methods.
While these reviews provide important insights into various fusion strategies, our work differs in its broader examination of technical challenges across diverse medical domains, along with a comprehensive exploration of solutions tailored to the clinical context. We aim to fill this gap by gathering the latest advancements in deep learning approaches for representing multimodal data and summarizing various research efforts in terms of the challenges they confront and the techniques they employ to overcome them. Although this systematic review focuses specifically on the challenges associated with multimodal data in the medical domains, some references included in the manuscript are not part of the systematic review. These additional references have been included to provide essential background information. By incorporating these sources, we aim to offer a more comprehensive context that complements the findings of the review.
@@ 2 Results
Figure 2: PRISMA flow diagram illustrating the study selection process following the systematic search.
@@
Intermediate fusion emerged as the most widely used approach, applied in 24 studies. Hybrid fusion was employed in 15 studies, while late fusion and early fusion strategies were utilized in 14 and 7 studies, respectively. Two studies employed hierarchical fusion strategies to tackle challenges in multimodal data integration 13, 14. Zhang et al. 14 introduced a transformer-based model that utilizes hierarchical multimodal transformers to progressively integrate features from shallow to deep layers. This design facilitates effective cross-modality interaction and addresses issues such as spatial misalignment and data heterogeneity, distinguishing it from traditional early, intermediate, or late fusion techniques. Similarly, Li et al. 13 proposed a hierarchical factorized bilinear fusion strategy tailored for cancer survival prediction, which integrates genomic and pathological image data across low- and high-level fusion stages. While both approaches highlight the critical role of cross-modality interactions, this method further incorporates modality-specific and cross-modality attentional factorized bilinear modules, enabling reduced computational complexity and enhanced performance, particularly in genomic data analysis. Together, these studies underscore the adaptability and effectiveness of hierarchical fusion frameworks in optimizing multimodal data integration across diverse applications. One study 15 employed a probability-based fusion approach, which involves calculating the probability of each feature being relevant and discriminative for classification, as determined by two optimization algorithms. Features with higher probabilities are prioritized and included in the final fused feature set, ensuring the combination of only the most informative and non-redundant features for classification. One other study 16 used the maximum correlation-based fusion approach, which emphasizes integrating features with the strongest cross-modal correlations. In this method, features from different modalities are analyzed for their pairwise correlations. Strongly correlated features are selected and combined, while weakly correlated or redundant features are excluded. This strategy enhances the representational power of the fused data by focusing on meaningful relationships while reducing noise and redundancy. Finally, five studies did not specify or clearly define their fusion strategy.
Table 2 provides an overview of the number of publications across different medical domains, along with the distribution of modeling challenges within these domains. This classification highlights the primary areas of application for multimodal data modeling, with neurology emerging as the most frequently studied domain (22 studies), followed by oncology (20 studies), physiology (6 studies), and ophthalmology and psychiatry (3 studies each). Additional domains such as psychology, endocrinology, cardiology, orthopedics, pharmacology, and gastroenterology further demonstrate the adaptability of multimodal data modeling across various medical contexts. Additionally, three studies explored non-medical applications focusing on synthetic data. Notably, neurology and oncology also stand out as the domains facing the highest number of modeling challenges, particularly regarding missing modalities, small datasets, and the identification of optimal fusion techniques.
The yearly frequencies of the identified challenges are illustrated in Figure Figure 3. Specifically since 2020, there was a strong increase in studies addressing challenges in modeling multimodal data (the number of studies in 2023 is lower as the search strategy was limited until October 2023). This upward trend indicates that research in this area will likely continue to expand, addressing new challenges as multimodal applications become more widespread.
@@
Medical Domain Number of Missing Interpretability Small Imbalance Optimal Other
Studies Data Data Dimensionality Fusion
@@
LABEL:Tab:challenges highlights for each study the particular challenges they tackled, along with the strategies they implemented to address these effectively. They are discussed in more detail and according to the challenges in the following section. Further explanation of the proposed solutions for each challenge, including the medical domain, disease type and fusion strategy used in each publication, can be found in the Supplementary file.
@@
Xing et al. 32   Gradient-guided knowledge modulation scheme
Zhang et al. 14   Throughout fusion transformer
Aghili et al. 33   Gradient boosting
Cai et al. 36   Vision Transformer
Hatami et al. 38   Weighted Ensenble Fusion
Karaman et al. 39   Re-weighted loss function
Lim et al. 41   Transfer learning
@@
Guan et al. 56   Knowledge distillation
Kustowski et al. 57   Transfer learning
Shickel et al. 59   Transfer learning
Wang et al. 67   Knowledge distillation
Tiulpin et al. 69   Using Gradient-weighted Class Activation Mapping
Cai et al. 71   Multimodal generative adversarial networks
Vivar et al. 72   Matrix completion
@@
Lei et al. 74   Dimension reduction/Feature selection
Eshaghi et al. 76   Kernel-Based   Heterogeneity: Kernel-Based
Thung et al. 79   Low-rank matrix completion algorithm
Tiwari et al. 81   Semi-Supervised Multi-Kernel Graph Embedding

### v3

# Challenges and proposed solutions in modeling multimodal data: A systematic review

@@ Abstract
In healthcare, data comes in various modalities like electronic health records (EHR), imaging, genomics, and wearable devices. Integrating more than one of these modalities in a multimodal data analysis holds significant promise for improving our understanding and management of complex health conditions 6. One notable domain where multimodal data analysis can make a transformative impact is in the management of chronic diseases such as asthma and chronic obstructive pulmonary disease (COPD) 2. Here, medical imaging data, such as computerized tomography, provides insights into lung morphology and pathological changes. Wearable devices monitor physiological parameters and activity levels, offering real-time data on the patient’s disease trajectory. Environmental data, encompassing air quality and pollutants, helps identify external factors influencing respiratory conditions. Clinical records and genomic information contribute to personalized treatment strategies, while patient-reported outcomes capture subjective experiences, shaping a patient-centric approach. Combining several or even all of these modalities allows us to understand the disease and reveal interactions between different aspects.
@@
Modeling multimodal data is a challenging task that extends well beyond the analysis of individual modalities. Specifically, each modality may have different data types, formats, scales, levels of noise, and levels of interpretability. Moreover, modalities may exhibit varying degrees of completeness, sparsity, and dimensionality, making it challenging to integrate and analyze them together. In the context of modeling multimodal data, fusion techniques are of great importance. A fusion strategy involves combining and integrating the different modalities to create a unified representation that captures the complementary information from each modality. These techniques can be divided into four main categories: (i) early fusion, (ii) intermediate fusion, (iii) late fusion, and (iv) hybrid fusion. (i) Early fusion involves integrating information from different sources at the data level without learning any marginal representations or feature extraction, treating them as unimodal input. Early fusion offers simplicity by integrating modalities at the data level, however, it often assumes that all input modalities are directly compatible and equally informative. It can capture basic cross-modal relationships present in raw data, but it typically struggles to model more complex, high-level interactions between modalities. In particular, early fusion may be inadequate when modalities differ in dimensionality, sampling rates, or semantic content, such as combining image data with structured omics data, where aligning and weighting heterogeneous features is non-trivial. (ii) Intermediate fusion, in contrast, involves learning modality-specific representations to identify correlations within each modality before generating joint representations or making predictions. This approach offers flexibility in selecting different architectures for feature extraction from various modalities 3. By employing distinct architectures and layers for each modality, it can effectively handle heterogeneous data. Additionally, intermediate fusion is resilient to dimensionality imbalances 3 and can accommodate missing modalities without significantly affecting performance 3, 4. (iii) Late fusion, on the other hand, trains separate models for each modality and integrates their individual predictions at a later stage. This method is particularly well-suited for managing heterogeneous modalities and addressing dimensionality imbalances and missing data. However, it may be less effective for modalities with strong correlations, as it does not capture interactions between different data sources. In addition, in scenarios where specific information holds greater significance, using late fusion offers notable benefits. The model prioritizes the most important information it has, then adds more details from other sources to improve its understanding. This prioritization allows for a more nuanced integration of information, enhancing the model’s ability to make informed decisions and predictions. (iv) Hybrid fusion combines elements from the aforementioned methods, allowing for flexible processing across multiple stages. This approach is especially effective in scenarios with diverse data modalities, where early fusion can be used for closely related data types, and later fusion stages facilitate the integration of more distinct modalities. It provides the adaptability needed to meet the specific requirements of multimodal models. The integration of diverse data modalities into a multimodal model is visually represented in Figure 1. The figure displays the four primary fusion strategies (early, intermediate, late, and hybrid), while other less frequently used strategies, such as hierarchical fusion, probability-based fusion, and maximum correlation-based fusion, are not shown here for simplicity but are briefly discussed in Section 2.
@@
Selecting the optimal fusion technique is only one of several challenges in multimodal modeling 5. This systematic review aims to identify the broader set of challenges encountered in modeling multimodal data for medical applications and to synthesize the solutions proposed in the reviewed studies. By structuring the review around key challenges such as missing modalities, small datasets, dimensionality imbalance, interpretability, and fusion strategies, we provide a challenge-oriented perspective that highlights not only the methodological progress achieved but also the gaps that remain.
@@
Several review articles have discussed the application of deep learning to analyze multimodal data. For instance, Mohsen et al. 6 explored integrating multiple data modalities with artificial intelligence models for a variety of applications, including combining medical imaging with data from electronic health records (EHRs). Similarly, Azam et al. 7 and Zhang et al. 8 solely focused on fusing different imaging modalities. Behrad and Abadeh 9 and Stahlschmidt et al. 3 primarily concentrated on integrating omics data with other modalities. Huang et al. 10 address the fusion of multimodal EHR data using traditional machine learning and deep learning methods. Pawłowski et al. 5 and Shaik et al. 12 focus on fusion techniques and feature extraction methods.
While these works provide valuable insights, e.g., into specific fusion techniques or modality combinations, none has systematically catalogued methods by the specific challenge they tackle across diverse medical applications. Our review addresses this gap by broadly examining technical challenges and compiling the solutions proposed for each, thus offering a challenge-oriented perspective that is currently missing from the literature. We aim to fill this gap by gathering the latest advancements in deep learning approaches for representing multimodal data and summarizing various research efforts in terms of the challenges they confront and the techniques they employ to overcome them.
The remainder of this paper is organized as follows: In Section 1, we present the methodology used for this systematic review, including the search strategy, inclusion and exclusion criteria, and data extraction process. Section 2 provides an overview of the challenges encountered in modeling multimodal data, including missing modalities, small sample sizes, dimensionality imbalance, interpretability, and identifying optimal fusion strategies. For each challenge, we summarize its impact on multimodal modeling and describe existing solutions reported in the reviewed studies. Section 3 discusses the key findings, highlights gaps in the literature, and reflects on limitations and future research directions. Finally, the paper concludes with Section 4, summarizing the main insights and contributions of this review.
1 Methods
1.1 Search strategy
To systematically identify relevant articles, a structured search strategy was employed (see Table 5). The search was conducted using two major databases, PubMed and manual search from Google Scholar. The time frame for the search was limited until October 2023 for PubMed (search date: 2023/10/09) and until end of 2023 for Google Scholar (search date: 2024/01/02). The strategy was designed to identify studies addressing challenges and methodologies in modeling multimodal data, with a focus on clinical applications. First, studies were required to explicitly mention “model fusion”, “data fusion,” or “multimodal” in their title, ensuring relevance to the topic of multimodal data modeling. To ensure comprehensive coverage, the Boolean operator OR was used within each keyword group to combine related terms. Next, the search included terms related to key research purposes, such as “disease prediction”, “diagnosis”, “clinical outcome prediction”, “modeling”, or “exacerbation”, to capture studies with clinical applications. Additionally, the search incorporated terms representing diverse data modalities, including “electronic health records”, “imaging data”, “wearable devices”, “longitudinal data”, “temporal data”, or “time-series data”, to ensure coverage of a wide range of multimodal datasets. To further refine the results, the term “challenges” was included to focus on studies discussing methodological or analytical obstacles related to multimodal data integration. Finally, the Boolean operator AND was used to combine these components. The final search query combined these elements to ensure a comprehensive yet focused selection of studies. In addition, a manual search was performed to identify any potentially relevant studies not captured in the database search. The reporting of this review follows the PRISMA 2020 guidelines page2021prisma.
1.2 Inclusion and exclusion criteria
1.3 Data extraction
For the data extraction process, we categorized the included publications according to the specific challenges they addressed. We identified five main technical challenges in modeling multimodal data: (i) missing entire modalities and incomplete datasets, (ii) small data (data sparsity), (iii) interpretability, (iv) imbalance in dimensionality, and (v) identifying the optimal fusion technique. We extracted information on the specific challenges each study tackled within the broader scope of the five identified challenges. We also captured the proposed solutions for overcoming these challenges. We further specifically collected information about the fusion strategies employed in each study. The choice of fusion technique directly relates to the challenge of finding the optimal fusion method (challenge v). Some strategies may also help mitigate other challenges, such as handling incomplete datasets (challenge i) or balancing dimensionality (challenge iv) by integrating information from multiple sources effectively.
To better understand which data types are most frequently integrated, we extracted the modality combinations reported in each study. We also recorded the type of analysis task (e.g., classification, clustering, survival prediction, …), as this directly determines the methodological setting in which multimodal data are analyzed. In addition, we noted whether studies reported quantitative performance metrics or comparative benchmarks against baseline methods. Capturing this information (e.g., accuracy, area under the curve (AUC), concordance index) was important because consistent reporting of evaluation measures is essential for assessing methodological impact and enabling meaningful comparisons across studies. Finally, because different medical settings encounter distinct modeling challenges, we recorded the clinical domain of each study to link it with the challenges addressed.
The screening process was conducted by five co-authors. The selected full-text studies were strategically distributed among three parties to ensure that each study was independently reviewed twice. Specifically, the first party (MF) screened the first two-thirds, the second party (MB, CG) handled the last two-thirds, and the third party (JK) examined the first and last thirds. In cases of discrepancies, an independent third screener (MW) was brought in to make a final decision.
@@ 2 Results
Figure 2: PRISMA flow diagram illustrating the study selection process following the systematic search in PubMed. In addition, 11 relevant studies were identified through a manual Google Scholar search, resulting in a total of 69 included studies.
@@
Intermediate fusion emerged as the most widely used approach, applied in 24 studies. Hybrid fusion was employed in 15 studies, while late fusion and early fusion strategies were utilized in 14 and 7 studies, respectively. Two studies employed hierarchical fusion strategies to tackle challenges in multimodal data integration One study introduced a transformer-based model that utilizes hierarchical multimodal transformers to progressively integrate features from shallow to deep layers. This design facilitates effective cross-modality interaction and addresses issues such as spatial misalignment and data heterogeneity, distinguishing it from traditional early, intermediate, or late fusion techniques. Similarly, one study proposed a hierarchical factorized bilinear fusion strategy tailored for cancer survival prediction, which integrates genomic and pathological image data across low- and high-level fusion stages. While both approaches highlight the critical role of cross-modality interactions, this method further incorporates modality-specific and cross-modality attentional factorized bilinear modules, enabling reduced computational complexity and enhanced performance, particularly in genomic data analysis. Together, these studies underscore the adaptability and effectiveness of hierarchical fusion frameworks in optimizing multimodal data integration across diverse applications. One study employed a probability-based fusion approach, which involves calculating the probability of each feature being relevant and discriminative for classification, as determined by two optimization algorithms. Features with higher probabilities are prioritized and included in the final fused feature set, ensuring the combination of only the most informative and non-redundant features for classification. One other study used the maximum correlation-based fusion approach, which emphasizes integrating features with the strongest cross-modal correlations. In this method, features from different modalities are analyzed for their pairwise correlations. Strongly correlated features are selected and combined, while weakly correlated or redundant features are excluded. This strategy enhances the representational power of the fused data by focusing on meaningful relationships while reducing noise and redundancy. Finally, five studies did not specify or clearly define their fusion strategy.
@@
Medical Domain
Number of
studies
Missing
data
Interpretability
Small
data
Imbalance
dimensionality
Optimal
fusion
Other
@@
Table 2 provides an overview of the number of publications across different medical domains, along with the distribution of modeling challenges within these domains. This classification highlights the primary areas of application for multimodal data modeling, with neurology emerging as the most frequently studied domain (22 studies), followed by oncology (20 studies), physiology (6 studies), and ophthalmology and psychiatry (3 studies each). Additional domains such as psychology, endocrinology, cardiology, orthopedics, pharmacology, and gastroenterology further demonstrate the adaptability of multimodal data modeling across various medical contexts. Additionally, three studies explored non-medical applications focusing on synthetic data. Notably, neurology and oncology stand out as the domains most affected by modeling challenges, where missing modalities and small datasets are especially frequent due to the difficulty of integrating heterogeneous data sources (e.g., imaging, genomics, and clinical notes) and the limited availability of samples for rare or complex conditions.
The yearly frequencies of the identified challenges are illustrated in Figure Figure 3. Specifically, since 2020, there was a strong increase in studies addressing challenges in modeling multimodal data (the number of studies in 2023 is lower as the search strategy was limited until October 2023). This upward trend indicates that research in this area will likely continue to expand, addressing new challenges as multimodal applications become more widespread.
LABEL:Tab:challenges highlights for each study the particular challenges they tackled, along with the strategies they implemented to address these effectively. They are discussed in more detail and according to the challenges in the following section. Further explanation of the proposed solutions for each challenge, including the medical domain, disease type, and fusion strategy used in each publication, can be found in Supplementary Table S1 for detailed per-study entries.
Following the overview of challenges and solutions in LABEL:Tab:challenges, we next analyzed which combinations of data modalities were most frequently investigated and how they relate to the type of task addressed. Table 4 and Figure 4 summarize these distributions. The majority of studies combined imaging with clinical data (22 studies) or imaging with genomic data (19 studies), reflecting their central role in current multimodal applications. These combinations were most often applied to classification tasks, with 20 studies using imaging + clinical data and 4 studies using imaging + genomic data, contributing to a total of 27 classification studies across all modality categories. In contrast, other modality pairings were far less common: for example, wearables combined with clinical data appeared in 5 studies, while multi-triple combinations such as imaging + genomic + clinical were reported in only 3 studies. Reported performance metrics also varied widely across tasks, with accuracy and AUC dominating classification studies (e.g., AUC values typically ranged from 0.85 to 0.95), while the concordance index (C-index) was most common for survival prediction (generally between 0.70 and 0.77). This heterogeneity underscores the lack of standardization in evaluation practices and complicates direct comparisons across studies. Overall, these distributions highlight both the dominant trends in multimodal modeling and potential opportunities for future research on underrepresented modality pairings and on harmonized performance reporting.
@@
Eshaghi et al. 76   Kernel-Based   Heterogeneity: Kernel-Based
Guan et al. 56   Knowledge distillation
Hatami et al. 38   Weighted Ensemble Fusion
@@
Kustowski et al. 57   Transfer learning
Lei et al. 74   Dimension reduction/Feature selection
@@
Thung et al. 79   Low-rank matrix completion algorithm
Tiulpin et al. 69   Using Gradient-weighted Class Activation Mapping
Tiwari et al. 81   Semi-Supervised Multi-Kernel Graph Embedding
Vivar et al. 72   Matrix completion
@@
Wang et al. 67   Knowledge distillation
Xing et al. 32   Gradient-guided knowledge modulation scheme
@@
Figure 4: Distribution of analysis tasks across data modality combinations. Bars show the number of studies per data modality, subdivided by task type.
Table 4: Studies grouped by data modalities, task type, and summarized performance metrics.
Data Modality Task (metric) + Studies Performance Metrics
Imaging + Genomic
Classification (AUC / Acc) 71, 17, 80, 26 AUC $\approx$ 0.93 71, 17; Acc= 0.90–0.92 80, 26
Survival prediction (C-index / AUC) 13, 40, 24 27, 28 C-index= 0.72–0.78 24, 27; AUC= 0.71–0.77 28; C-index= 0.766 13, 40
Prediction 21, 25, 43, 45 AUC= 0.85 43; AUC= 0.80 45;
AUROC= 0.898 21; AUROC= 0.806 25
Joint (Class. + Regr.) 52, 62 RMSE= 10 52; RMSE= 16.2 ± 0.76 62
Representation learning 61 Acc= 0.90
Regression (RMSE) 73 Not reported
Detection 70 $p$-value=0.000001
Imaging + Clinical
Classification (AUC / Acc) 33, 36, 19, 55 76, 56, 15, 16, 74, 22, 41, 65 77, 64, 58, 23, 44, 81 72, 46, 30, 60, xu_explainable_2021, 14 not reported 33, 22; Acc=97% 19; Acc= 94% 55; Acc= 88% 76; Acc= 0.80 56; Acc=99.1% 15; Acc=99.46% 16; Acc= 97% 74; AUC= 0.86 65; Acc= 91% 77; C-index= 0.86–0.94
64; Acc= 93% 58; Acc= 94% 23; Acc= 75% 44; AUC= 0.90 81; AUC= 0.95 72; AUC= 0.91 xu_explainable_2021; AUC= 0.94 41; Acc= 0.91 46; AUC= 0.9 30
Prediction 38, 39, 57, 63, 59, 69, 47, 49 $wR\approx 0.72\text{--}0.80$ 49; AUROC= 0.92 59;
Not reported 57; AUROC= 0.90 63;
AUC= 0.77 38; ROC $\approx$ 0.90 39; AUC= 0.79 69;
MAPE= 0.165 47
Survival prediction (C-index) 42, 50 AUC= 0.75 42; Acc= 88% 50
Joint (Class. + Regr.) 79 Acc= 0.912 79
Generation / Reconstruction 51 Not reported
Genomic + Clinical
Survival prediction (C-index) 68, 20, C-index= 0.63 68; C-index= 0.60–0.75 20
Other (Federated learning) 18 Not reported
Imaging + Genomic + Clinical
Classification (AUC / Acc) 37, 32, Acc= 96.9% 37, AUC= 0.92 32
Biomarker discovery / diagnostics 35 Not reported
Wearable + Clinical
Classification (Acc) 58 Acc= 93.4% 58
Representation / Feature learning 54, 31 Acc= 95% 54; Acc= 59% 31
Joint (Class. + Regr.) 29 AUPRC= 0.56 29
Regression (MAE) 66 RMSE reduction=55% 66
Other multimodal (synthetic / simulation / non-medical)
Classification (Acc) 67 Acc= 75%
Representation learning (Acc) 78 Acc= 95% (ETH-80);
Prediction (Bias correction) 57 not reported in standard biomedical metrics
Authentication (Error rates) 75 reduced EER by up to 32% and HTER by 28%
Clinical / EHR
Prediction (AUROC) 34, 53, 63 AUROC = 0.84–0.96 34; AUC= 0.94;
F1= 0.88 53, AUROC= 0.85–0.90 63